Shift summary for eve. 10/6/99

Ed Blucher (blucher@hep.uchicago.edu)
Wed, 6 Oct 1999 23:44:59 -0500


Shift Summary for Eve Shift, 10/6/99

On Shift: Patty McBride, Ed Blucher

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

========================================================================
Overview of shift:
No beam. TeV purification in progress. Nothing going on at KTeV.

========================================================================
Description of Shift:

16:35 Dave N. installs two new argon-ethane bottles on center island.

17:45 Angela takes DC gas samples.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift